OpenStudy (anonymous):

-7x - 4y = 23 2x + 8y = 14 solve by elimination.

OpenStudy (bahrom7893):

2x + 8y = 14 x+4y=7 x = 7 - 4y

OpenStudy (bahrom7893):

plug that into the first equation: -7(7-4y)-4y=23 -49+28y-4y=23

OpenStudy (anonymous):

why do u have x + 4y = 7

OpenStudy (bahrom7893):

Divide everything in the 2nd equation by 2

OpenStudy (anonymous):

and then after u plug it, that's when i get confused

OpenStudy (bahrom7893):

-7(7-4y)-4y=23 -7*7+(-7)*(-4y)-4y=23

OpenStudy (bahrom7893):

-49 + 28y - 4y = 23

OpenStudy (anonymous):

hold up, i think i got it:) thank you! i have 2 mrore questions, if you have time.

OpenStudy (anonymous):

more*

OpenStudy (anonymous):

24y = 12x - 24 8x = 16y + 16

OpenStudy (bahrom7893):

look at the 2nd equation, divide everything by 8: x = 2y+2

OpenStudy (anonymous):

\[-7x-4y=23\] \[2x+8y=14\] We should be able to get at least one number common in both equations to eliminate it, so multiply 1st equation by -2 to get 8y in both equation one and two, so we have: \[14x+8y=-46\] \[2x+8y=14\] Cancel out 8y so, we have; \[14x-2x=-46-14\] \[12x=-60\] \[x=-60/12\] \[x=-5\] Substitute the value of x in any of the two equations to find y: \[2(-5)+8y=14\] 8y=14+10 8y=24 y=24/8=3 y=3 So, x=-5 and y=3

OpenStudy (anonymous):

okay, and what should i do next? :)

OpenStudy (bahrom7893):

24y = 12x - 24 8x = 16y + 16 look at the 2nd equation, divide everything by 8: x = 2y+2 Divide the first equation by 12: 2y = x - 2 Plug the value of x from the 2nd equation into the first one: 2y = 2y+2-2 0=0 Those are both the same line, an infinite number of solutions exists

OpenStudy (anonymous):

go go go bahrom! lol

OpenStudy (anonymous):

oh okayy! :) & last one; 0 = -24 - 24y + 16x 2 - y = -x

OpenStudy (bahrom7893):

divide the top one by 4: 0 = -6 - 6y +4x multiply everything by -1 in the 2nd equation: -2+y=x, or x = y-2 Plug that into the first equation: 0 = -6 - 6y +4x 0 = -6 - 6y +4(y-2) 0 = -6 - 6y +4y-8 0 = -6-8-6y+4y 0=-14-2y add 14 to both sides: 14 = -2y divide both sides by -2: 14/-2 = y -7=y, or y = -7

OpenStudy (bahrom7893):

x = y-2 x = -7-2=-9

OpenStudy (anonymous):

thank you!!!!